### v1

@@ Abstract
When aggregating Subjective Expected Utility preferences, the Pareto principle leads to an impossibility result unless the individuals have a common belief. This paper examines the source of this impossibility in more detail by considering the aggregation of a general class of incomplete preferences that can represent gradual ambiguity perceptions. Our result shows that the planner cannot avoid ignoring some individuals unless there is a probability distribution that all individuals agree is most plausible. This means that even if individuals have similar ambiguity perceptions, the impossibility persists as long as some individual’s most plausible belief differs even slightly from those of others.
@@ 1 Introduction
However, it is known that this principle leads to unacceptable conclusions under uncertainty. Hylland and Zeckhauser (1979) and Mongin (1995, 1998) showed that when aggregating Subjective Expected Utility preferences, the Pareto principle leads to dictatorships if there is a disagreement of beliefs. Following them, many papers have examined the preference aggregation under uncertainty (e.g., Gilboa et al. (2004); Chambers and Hayashi (2006); Alon and Gayer (2016); Danan et al. (2016); Zuber (2016); Billot and Qu (2021)).
@@
Another approach to relaxing the assumption of social rationality is to drop completeness or transitivity. Danan et al. (2016) studied the difficulty of the Pareto principle when aggregating Bewley preferences, an incomplete multiprior version of the Subjective Expected Utility model (cf. Bewley (2002)). They showed that the Pareto principle implies that (i) vNM functions are aggregated by a weighted utilitarian rule as in Harsanyi (1955) and (ii) the planner’s belief set is in the intersection of belief sets of individuals who are assigned positive weights. Therefore, if the intersection of all individuals’ belief sets is empty, then the planner has to ignore some individuals’ preferences by assigning them a weight of 0, that is, the impossibility still holds.
@@
Our main result shows that the Pareto principle leads to impossibility again when aggregating the above general preferences, and clarifies the conditions under which this principle leads to negative consequences. Roughly speaking, the planner cannot avoid ignoring some individuals unless there is a probability distribution that all individuals unanimously consider most plausible. In other words, even if individuals have similar ambiguity perceptions, the impossibility holds as long as some individual’s most plausible belief differs even slightly from those of others. Therefore, our result shows that it is difficult to satisfy the Pareto principle and avoid ignoring some individuals at the same time, even if we consider a social preference that can respect individuals’ beliefs flexibly.
@@
Preference aggregation with gradual ambiguity perceptions was also considered in Dong-Xuan (2024) and Mudekereza (2025). They mainly considered the decision-making model examined in Cerreia-Vioglio et al. (2025), which is a special case of variational preferences (Maccheroni et al. (2006)). Dong-Xuan studied the case where all individuals have a common vNM function, so focused on the aggregation of ambiguity perceptions. Mudekereza examined the sequential process where the planner first aggregates the subrelations induced by the individual preferences and then selects the optimal probability model (i.e., a set of priors). Since these subrelations can be represented by the Bewley model, the aggregation step is essentially based on the theorems of Danan et al. (2016), as Mudekereza argued.
This paper is organized as follows: Section 2 introduces the setup studied in this paper. Section 3 gives the main results and two corollaries. Section 4 discusses our result in more detail.
@@ 3 Results
Before providing a proof, we discuss the implications of this result. This theorem states that vNM functions are aggregated in a utilitarian way as Harsanyi’s (1955) theorem. Furthermore, the planner’s perception function is bounded below using the individuals’ perception functions and the weight vector used to aggregate the vNM functions. To see this condition in more detail, we consider a result immediately following from Theorem 1.
@@
Compared to this result, our results show that the difficulty of Standard Pareto is more severe than expected. Corollary 1 states that if the intersection of sets of probability distributions each individual thinks to be most plausible is empty, then the planner cannot avoid excluding some individuals’ preferences. This means that even if the individuals have perception functions that are very similar to but slightly different from each other, then the planner cannot respect all individuals’ preferences and, moreover, has to evaluate all alternatives based on only one individual’s preference.
@@ 4 Discussion
Finally, we discuss the violation of transitivity in the social preference and the converse of Theorem 1.
@@ 4.1 Violation of Transitivity and Independence
As noted in Remark 1, variational Bewley preferences may violate transitivity and independence. While considering such individual preferences may be appropriate for describing actual individual choice patterns, it is more controversial whether it is appropriate to use such preferences for social evaluation. In this section, we discuss the necessity and validity of considering such preferences as those of the social planner.
First, violations of these axioms are needed to account for gradually changing perceptions of ambiguity. Starting from the Subjective Expected Utility model, the Bewley model can be characterized by the violation of completeness. Apart from transitivity and independence, other axioms in the Subjective Expected Utility model—nontriviality, monotonicity, and continuity—are either very basic or merely technical. Therefore, to accommodate gradual ambiguity perceptions, it is necessary to consider preferences that may violate transitivity and independence. Furthermore, it should be noted that dropping one of these two is insufficient: As discussed in Remark 1, a variational Bewley preference that satisfies transitivity or independence is a Bewley preference.
Second, if we allow individual preferences to violate transitivity and independence, a social preference derived from them can also easily violate these axioms. To illustrate this, consider an extreme example in which all individuals share a common variational Bewley preference that cannot be represented by the standard Bewley model. Since the social preference should be based on the individuals’ preferences, we would require the social ranking to completely align with individual rankings. However, in this case, the social preference cannot satisfy transitivity and independence because the individual preferences themselves do not satisfy these properties. Therefore, imposing these axioms solely on the social preference would exclude these intuitive aggregation rules.
Notice that the literature has also considered social preferences that may violate transitivity and independence. In the context of Arrovian preference aggregation, intransitive social preferences have been examined (e.g., Sen (1969); Mas-Colell and Sonnenschein (1972); Blau and Deb (1977); Blair and Pollak (1982); Brandl and Brandt (2020)). For independence, since Diamond’s (1967) critique of Harsanyi (1955), many papers have studied social preferences that are not assumed to satisfy independence (e.g., Alon and Gayer (2016); Hayashi and Lombardi (2019); Mongin and Pivato (2021)).
However, of course, we can take the position that the social preference should satisfy transitivity and independence. Since the social preference becomes a Bewley preference in this case, the following result can be obtained from our main result:
@@ 4.2 The Converse of Theorem 1
Recall that Theorem 1 is not an if-and-only-if statement because the converse does not hold. To see this, consider a two-person and two-state example.
@@
Let $N=\{1,2\}$ and $S=\{s_{0},s_{1}\}$. In this case, each probability distribution over $S$ can be identified with a probability $p_{1}$ of $s_{1}$. Consider two acts $f$ and $g$ and the following preference profile: $u_{1}(f(s_{0}))=u_{1}(g(s_{0}))=u_{2}(f(s_{0}))=u_{2}(g(s_{0}))=0$, $u_{1}(f(s_{1}))=u_{2}(f(s_{1}))=2$, $u_{1}(g(s_{1}))=u_{2}(g(s_{1}))=4$, and $c_{1}(p_{1})=c_{2}(p_{1})=2p_{1}$ for each $p_{1}\in[0,1]$. Then, $f\sim_{1}g$ and $f\sim_{2}g$ hold. Consider the aggregation rule such that $u_{0}={1\over 2}u_{1}+{1\over 2}u_{2}$ and $c_{0}={1\over 2}c_{1}$. Then, this social preference violates Standard Pareto although it satisfies the conditions in Theorem 1.777Note that we can construct such a preference profile without violating Preference Diversity. For instance, suppose that $X=\mathbb{R}^{3}$ and that $u_{1}(x)=x_{1}+x_{3}$ and $u_{2}(x)=x_{2}+x_{3}$ for all $x\in X$. Then, Preference Diversity holds, and we can take $f,g\in F$ satisfying the desired conditions by choosing $f(s)$ and $g(s)$ from $\{(0,0,x_{3})\mid x_{3}\in\mathbb{R}\}$ for each $s\in S$. Indeed, if $p_{1}=1$, then $u_{0}(f(s_{1}))+c_{0}(1)={1\over 2}u_{1}(f(s_{1}))+{1\over 2}c_{1}(1)+{1\over 2}u_{2}(f(s_{1}))=3<4={1\over 2}u_{1}(g(s_{1}))+{1\over 2}u_{2}(g(s_{1}))=u_{0}(g(s_{1}))$, which implies $f\not\succsim_{0}g$ (i.e., $g\succ_{0}f$).
Thus, our result provides only weak conditions obtained from Standard Pareto. Since we understand Theorem 1 as an impossibility result, this problem is not so severe. However, it is still important to examine when the converse holds, because this helps clarify the conditions under which the social planner’s perception function can be constructed without violating the imposed axiom. To address this issue, we here discuss (i) sufficient conditions under which the converse holds and (ii) an axiom that characterizes the aggregation rule obtained in Theorem 1.
First, we consider the issue (i). We can show that if the preferences of all individuals or the social planner’s preferences are assumed to be Bewley preferences, then Theorem 1 becomes the if-and-only-if result. In other words, to satisfy Standard Pareto, it is sufficient that either the planner or the individuals have preferences with transitivity and independence (cf. Remark 1 (iv)). For the case where the planner’s preference is a Bewley preference, see Proposition 1. The following result considers the case where individual preferences are Bewley preferences.
@@
Next, we consider the issue (ii), that is, the axiom that is equivalent to the aggregation rule we have obtained in Theorem 1. For $i\in N$, let $\mathcal{F}^{i}\subseteq\mathcal{F}$ be such that if $f\in\mathcal{F}^{i}$, then for each $s\in S$, $f(s)$ is on the line passing through $\bar{x}^{i}$ and $\underline{x}^{i}$, where these two are the elements in $X$ introduced in the definition of Preference Diversity. For all $f,g\in\mathcal{F}^{i}$, by the definition of $\bar{x}^{i}$ and $\underline{x}^{i}$, we have $f\sim_{j}g$ for all $j\in N\setminus\{i\}$. Therefore, $\mathcal{F}^{i}$ is the set of acts that is relevant only for individual $i$.
We then consider an axiom much weaker than Standard Pareto.
@@
This axiom postulates that for any pair of alternatives that is relevant only for individual $i$ (e.g., the color of shirts that $i$ wears), the planner should respect $i$’s opinion (cf. Sen (1970)). The following result shows that Liberalism is a necessary and sufficient condition for the aggregation rules obtained in Theorem 1.
@@
Suppose that for each $\in N\cup\{0\}, i\succsim_{i}$ is a variational Bewley preference $(u_{i},c_{i})$ and Preference Diversity holds. Then, Liberalism holds if and only if there exists $(\alpha,\beta)\in\mathbb{R}^{n}_{+}\times\mathbb{R}$ such that
@@
We only prove the if part since the only-if part can be shown in the same way as Theorem 1. Suppose that for some $(\alpha,\beta)\in\mathbb{R}^{n}_{+}\times\mathbb{R}$, (8) holds. Take $i\in N$ arbitrarily. Let $f,g\in\mathcal{F}^{i}$ such that $f\succsim_{i}g$, that is, $\sum_{s\in S}p(s)u_{i}(f(s))+c_{i}(p)\geq\sum_{s\in S}p(s)u_{i}(g(s))$ for all $p\in\Delta(S)$. By the definition of $\mathcal{F}^{i}$, $u_{j}(f(s))=u_{j}(g(s))$ for all $j\in N\setminus\{i\}$ and $s\in S$. Therefore, by (8),
@@
Therefore, Liberalism holds. ∎
Although such an if-and-only-if statement holds, we have used Standard Pareto instead of Liberalism in the main part of this paper because Standard Pareto is a simpler and more basic condition, which can be defined without depending on Preference Diversity. Moreover, starting with Standard Pareto motivates us to conduct the subsequent analysis, such as Propositions 1 and 2.

### v2

@@ Abstract
When aggregating Subjective Expected Utility preferences, the Pareto principle leads to an impossibility result unless the individuals have a common belief. This paper examines the source of this impossibility in more detail by considering the aggregation of a general class of incomplete preferences that can represent gradual ambiguity perceptions. Our result shows that the planner cannot avoid ignoring some individuals unless there is a probability distribution that all individuals agree is most plausible. This means that even if individuals have similar ambiguity perceptions, the impossibility persists as long as some individual’s most plausible belief differs even slightly from that of others.
@@ 1 Introduction
However, it is known that this principle leads to unacceptable conclusions under uncertainty. Hylland and Zeckhauser (1979) and Mongin (1995, 1998) showed that when aggregating Subjective Expected Utility preferences, the Pareto principle leads to dictatorships if there is a disagreement of beliefs. Following them, many papers have examined preference aggregation under uncertainty (e.g., Gilboa et al. (2004); Chambers and Hayashi (2006); Alon and Gayer (2016); Danan et al. (2016); Zuber (2016); Billot and Qu (2021)).
@@
Another approach to relaxing the assumption of social rationality is to drop completeness or transitivity. Danan et al. (2016) studied the difficulty of the Pareto principle when aggregating Bewley preferences, an incomplete multiprior version of the Subjective Expected Utility model (cf. Bewley (2002)). They showed that the Pareto principle implies that (i) vNM functions are aggregated by a weighted utilitarian rule à la Harsanyi (1955) and (ii) the planner’s belief set is in the intersection of belief sets of individuals who are assigned positive weights. Therefore, if the intersection of all individuals’ belief sets is empty, then the planner has to ignore some individuals’ preferences by assigning them a weight of 0, that is, the impossibility still holds. [^2] See also Pivato and Tchouante (2024) and Kurata and Nakamura (2025) for the aggregation of Bewley preferences.
@@
Our main result shows that the Pareto principle leads to impossibility again when aggregating the above general preferences, and clarifies a sufficient condition under which this principle leads to negative consequences. Roughly speaking, the planner cannot avoid ignoring some individuals unless there is a probability distribution that all individuals unanimously consider most plausible. In other words, even if individuals have similar ambiguity perceptions, the impossibility holds as long as some individual’s most plausible belief differs even slightly from those of others. Therefore, our result shows that it is difficult to satisfy the Pareto principle and avoid ignoring some individuals at the same time, even if we consider a social preference that can respect individuals’ beliefs flexibly.
@@
Preference aggregation with gradual ambiguity perceptions was also considered in Dong-Xuan (2024) and Mudekereza (2025). They mainly considered the decision-making model examined in Cerreia–Vioglio et al. (2025), which is a special case of variational preferences (Maccheroni et al. (2006)). Dong-Xuan studied the case where all individuals have a common vNM function and thus focused on the aggregation of ambiguity perceptions. Mudekereza examined the sequential process where the planner first aggregates the subrelations induced by the individual preferences and then selects the optimal probability model (i.e., a set of priors). Since these subrelations can be represented by the Bewley model, the aggregation step is essentially based on the theorems of Danan et al. (2016), as Mudekereza argued.
This paper is organized as follows: Section 2 introduces the setup studied in this paper. Section 3 gives the main results and two corollaries. Section 4 discusses our result in more detail. Section 5 provides concluding remarks.
@@ 3 Results
Before providing a proof, we discuss the implications of this result. This theorem states that vNM functions are aggregated in a utilitarian way, as in Harsanyi’s (1955) theorem. Furthermore, the planner’s perception function is bounded below using the individuals’ perception functions and the weight vector used to aggregate the vNM functions. To see this condition in more detail, we consider a result immediately following from Theorem 1.
@@
Compared to this result, our results show that the difficulty of Standard Pareto is more severe than expected. Corollary 1 states that if the intersection of sets of probability distributions each individual thinks to be most plausible is empty, then the planner cannot avoid excluding some individuals’ preferences. This means that even if the individuals have perception functions that are very similar to but slightly different from one another, the planner cannot respect all individuals’ preferences and, moreover, has to evaluate all alternatives based on only one individual’s preference.
@@ 4 Discussion
Finally, we discuss the violations of transitivity and independence in the social preference and the converse of Theorem 1.
@@ 4.1 Violation of Transitivity and Independence
As noted in Remark 1, variational Bewley preferences may violate transitivity and independence. While considering such individual preferences may be appropriate for describing actual individual choice patterns, it is more controversial whether it is appropriate to use such preferences for social evaluation. [^8] Notice that the literature has also considered social preferences that may violate transitivity and independence. In the context of Arrovian preference aggregation, intransitive social preferences have been examined (e.g., Sen (1969); Mas-Colell and Sonnenschein (1972); Blau and Deb (1977); Blair and Pollak (1982); Brandl and Brandt (2020)). For independence, since Diamond’s (1967) critique of Harsanyi (1955), many papers have studied social preferences that are not assumed to satisfy independence (e.g., Alon and Gayer (2016); Hayashi and Lombardi (2019); Mongin and Pivato (2021)). We argue that to account for gradual ambiguity perceptions, studying preferences that violate these axioms would be one of the reasonable ways. Starting from the Subjective Expected Utility model, the Bewley model can be characterized by the violation of completeness. Apart from transitivity and independence, the remaining axioms in the Subjective Expected Utility model—nontriviality, monotonicity, and continuity—are either very basic or merely technical. Therefore, to accommodate gradual ambiguity perceptions, it is necessary to consider preferences that may violate transitivity or independence. Once the above argument is accepted, dropping one of these two is insufficient: As discussed in Remark 1, a variational Bewley preference that satisfies transitivity or independence is a Bewley preference.
On the other hand, of course, we can take the position that the social preference should satisfy transitivity and independence. Since the social preference becomes a Bewley preference in this case, the following result can be obtained from our main result:
@@ 4.2 The Converse of Theorem 1
Recall that Theorem 1 is not an if-and-only-if statement because the converse does not hold. To see this, consider a two-person, two-state example.
@@
Let $N=\{1,2\}$ and $S=\{s_{0},s_{1}\}$. In this case, each probability distribution over $S$ can be identified with a probability $p_{1}$ of $s_{1}$. Consider two acts $f$ and $g$ and the following preference profile: $u_{1}(f(s_{0}))=u_{1}(g(s_{0}))=u_{2}(f(s_{0}))=u_{2}(g(s_{0}))=0$, $u_{1}(f(s_{1}))=u_{2}(f(s_{1}))=2$, $u_{1}(g(s_{1}))=u_{2}(g(s_{1}))=4$, and $c_{1}(p_{1})=c_{2}(p_{1})=2p_{1}$ for each $p_{1}\in[0,1]$. Then, $f\sim_{1}g$ and $f\sim_{2}g$ hold. Consider the aggregation rule such that $u_{0}={1\over 2}u_{1}+{1\over 2}u_{2}$ and $c_{0}={1\over 2}c_{1}$. Then, this social preference violates Standard Pareto although it satisfies the conditions in Theorem 1.999Note that we can construct such a preference profile without violating Preference Diversity. For instance, suppose that $X=\mathbb{R}^{3}$ and that $u_{1}(x)=x_{1}+x_{3}$ and $u_{2}(x)=x_{2}+x_{3}$ for all $x\in X$. Then, Preference Diversity holds, and we can take $f,g\in F$ satisfying the desired conditions by choosing $f(s)$ and $g(s)$ from $\{(0,0,x_{3})\mid x_{3}\in\mathbb{R}\}$ for each $s\in S$. Indeed, if $p_{1}=1$, then $u_{0}(f(s_{1}))+c_{0}(1)={1\over 2}u_{1}(f(s_{1}))+{1\over 2}c_{1}(1)+{1\over 2}u_{2}(f(s_{1}))=3<4={1\over 2}u_{1}(g(s_{1}))+{1\over 2}u_{2}(g(s_{1}))=u_{0}(g(s_{1}))$, which implies $f\not\succsim_{0}g$.
Thus, our result provides only weak conditions obtained from Standard Pareto. Since we understand Theorem 1 as an impossibility result, this problem is not so severe. However, it is still important to examine when the converse holds, because this helps clarify the conditions under which the social planner’s perception function can be constructed without violating the imposed axiom. To address this issue, we here discuss (i) sufficient conditions under which the converse holds and (ii) an axiom that characterizes the aggregation rules obtained in Theorem 1.
First, we consider issue (i). We can show that if the preferences of all individuals or the social planner’s preferences are assumed to be Bewley preferences, then Theorem 1 becomes an if-and-only-if result. In other words, to satisfy Standard Pareto, it is sufficient that either the planner or the individuals have preferences with transitivity and independence (cf. Remark 1 (iv)). For the case where the planner’s preference is a Bewley preference, see Proposition 1. The following result considers the case where individual preferences are Bewley preferences.
@@
Next, we address issue (ii), that is, axioms that characterize the aggregation rules in Theorem 1. For $i\in N$, let $\mathcal{F}^{i}\subseteq\mathcal{F}$ be such that if $f\in\mathcal{F}^{i}$, then for each $s\in S$, $f(s)$ is on the line passing through $\bar{x}^{i}$ and $\underline{x}^{i}$, where these two are the elements in $X$ introduced in the definition of Preference Diversity. For all $f,g\in\mathcal{F}^{i}$, by the definition of $\bar{x}^{i}$ and $\underline{x}^{i}$, we have $f\sim_{j}g$ for all $j\in N\setminus\{i\}$. Therefore, $\mathcal{F}^{i}$ is the set of acts that is relevant only for individual $i$.
We then consider three axioms weaker than Standard Pareto.
Constant Pareto.
For all $x,y\in X$, if $x\succsim_{i}y$ for all $i\in N$, then $x\succsim_{0}y$.
@@
Dominance Pareto.
For all $f,g\in\mathcal{F}$, if $f\succsim_{i}g$ for some $i\in N$ and $f(s)\succsim_{j}g(s)$ for all $s\in S$ and $j\in N\backslash\{i\}$, then $f\succsim_{0}g$.
Liberalism postulates that for any pair of alternatives that is relevant only for individual $i$ (e.g., the color of shirts that $i$ wears), the planner should respect $i$’s opinion (cf. Sen (1970)). Dominance Pareto considers the case where all individuals except for individual $i$ think an act $f$ is weakly better than another act $g$ in every state. Then, perception functions of these individuals do not matter, and hence only the perception function of $i$ is relevant. This axiom requires that if $i$ also weakly prefers $f$ to $g$ in such a situation, then $f$ should be socially at least as good as $g$. The following result characterizes the aggregation rules obtained in Theorem 1.101010The author is grateful to a referee for suggesting Dominance Pareto.
@@
Suppose that for each $\in N\cup\{0\}, i\succsim_{i}$ is a variational Bewley preference $(u_{i},c_{i})$ and Preference Diversity holds. Then, the following three statements are equivalent:
  1. (i)
    Constant Pareto and Liberalism hold;
  2. (ii)
    Dominance Pareto holds; and
  3. (iii)
    there exists $(\alpha,\beta)\in\mathbb{R}^{n}_{+}\times\mathbb{R}$ such that
@@
We can prove that each of (i) and (ii) implies (iii) in the same way as Theorem 1. We show that (iii) implies (i). Suppose that (8) holds for some $(\alpha,\beta)\in\mathbb{R}^{n}_{+}\times\mathbb{R}$. It is straightforward to check that Constant Pareto holds. Take $i\in N$ arbitrarily. Let $f,g\in\mathcal{F}^{i}$ such that $f\succsim_{i}g$, that is, $\sum_{s\in S}p(s)u_{i}(f(s))+c_{i}(p)\geq\sum_{s\in S}p(s)u_{i}(g(s))$ for all $p\in\Delta(S)$. By the definition of $\mathcal{F}^{i}$, $u_{j}(f(s))=u_{j}(g(s))$ for all $s\in S$ and $j\in N\setminus\{i\}$. Therefore, by (8),
@@
Therefore, Liberalism holds.
To prove that (iii) implies (ii), suppose that (iii) holds again. Let $f,g\in\mathcal{F}$ be such that $f\succsim_{i}g$ for some $i\in N$ and $f(s)\succsim_{j}g(s)$ for all $s\in S$ and $j\in N\backslash\{i\}$. Then, we have $\sum_{s\in S}p(s)u_{i}(f(s))+c_{i}(p)\geq\sum_{s\in S}p(s)u_{i}(g(s))$ and $u_{j}(f(s))\geq u_{j}(g(s))$ for all $s\in S$ and $j\in N\setminus\{i\}$. As in the last paragraph, we can show that $\sum_{s\in S}p(s)u_{0}(f(s))+c_{0}(p)\geq\sum_{s\in S}p(s)u_{0}(g(s))$ for all $p\in\Delta(S)$. Therefore, Dominance Pareto holds. ∎
Although such an if-and-only-if result holds, we have used Standard Pareto in the main part of this paper because Standard Pareto is a simpler and more basic condition, which can be defined without relying on Preference Diversity. Moreover, starting with Standard Pareto motivates us to conduct the subsequent analysis, such as Propositions 1 and 2.
5 Concluding Remarks
We have examined the Paretian aggregation of variational Bewley preferences and showed that under Standard Pareto, the social planner cannot respect all individual preferences in most cases. While Danan et al. (2016) obtained a possibility result using a restricted Pareto principle, whether we can obtain such a possibility result by weakening Standard Pareto in a reasonable way remains an open question.
We finally discuss that the approach taken by Danan et al. (2016) cannot work well in our framework. We say that two acts $f$ and $g$ are common-taste acts if $x\succsim_{i}y$ is equivalent to $x\succsim_{j}y$ for all $x$ and $y$ in the convex hull of $f(S)\cup g(S)$ and all $i,j\in N$. They examined the condition that requires that for all common-taste acts $f,g\in\mathcal{F}$, if $f\succsim_{i}g$ for all $i\in N$, then $f\succsim_{0}g$. By imposing the Paretian condition on pairs without taste disagreements, we can separate the perception aggregation from the taste aggregation.
To conduct a similar analysis, we need to introduce an additional ad-hoc assumption in the above condition to focus on a specific pair of outcomes: for instance, we need to consider the Paretian principles restricted to the pairs of acts on $\{f\in\mathcal{F}\mid\forall s\in S,~\exists\alpha\in[0,1]~~\text{s.t.}~~f(s)=\alpha x^{+}+(1-\alpha)x^{-}\}$ for some fixed $x^{+},x^{-}\in X$, and the obtained aggregation rules depend on the choice of $x^{+}$ and $x^{-}$. The key reason behind the difference from Danan et al. (2016) lies in the uniqueness of parameters in the variational Bewley model. In contrast to the Bewley model, the uniqueness of perception functions can be obtained only when we fix some vNM function in variational Bewley preferences. Due to this difference, we cannot deal with these parameters separately, which is allowed in the Bewley model. Therefore, even if we focus on acts with no disagreement in tastes, we cannot concentrate on the aggregation of perception functions.